OpenStudy (jordanloveangel):

Louann and Carla received equal scores on a test made up of multiple choice questions and an essay. Louann got 14 multiple choice questions correct and received 24 points for her essay. Carla got 16 multiple choice questions correct and received 14 points for her essay. How many points was each multiple choice question worth? ?points

OpenStudy (jordanloveangel):

@cody7744

OpenStudy (igreen):

It's 5..because: 5 * 14 = 70. 70 + 24 = 94. That means Louann's score is 94 if each Multiple choice question is worth 5 points. 16 * 5 = 80 80 + 14 = 94. That means Carla's score is 94 if each Multiple choice question is worth 5 points also.. Therefore, the answer is 5.

OpenStudy (jordanloveangel):

thanx